OpenStudy (anonymous):

4/2 and 16/12 do these pair of ratios form a proportion

OpenStudy (anonymous):

Not sure what you asking

OpenStudy (anonymous):

4/2=2 and 16/12= 4/3

OpenStudy (anonymous):

The ratios are not proportional to one another.

OpenStudy (anonymous):

Proportion to each other?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

what grade material is this?

OpenStudy (anonymous):

Just making sure we are on the same page

OpenStudy (anonymous):

oh ok ask away :)

OpenStudy (anonymous):

Try it yourself first

OpenStudy (anonymous):

What's 3/2?

OpenStudy (anonymous):

ok haveagood one